OpenStudy (anonymous):

how do i add exponents again is it like only like variables such as x^2+x^2 or can X^3+x^2 work? im thinking the second

OpenStudy (mertsj):

Yes. and only if you are multiplying. NOT for addition.

OpenStudy (mertsj):

x^2+x^2=2x^2 x^3 and x^2 are not like terms and cannot be added. x^3 times x^2=x^5

OpenStudy (marissalovescats):

^^ Yup. Like if you have 10^3*10^17 it would = 10^20

OpenStudy (anonymous):

so in addition unlike exponents is jsut left alone such as x^3+x^4+x^5?

OpenStudy (marissalovescats):

No you cannot add exponent powers. Only if you are multiplying something with the same base.

OpenStudy (marissalovescats):

Do you understand?

OpenStudy (anonymous):

yes sorry facebook lol